Brokeback Towing Company is at the end of its accounting year, December 31, 2018. The following data that must be considered were developed from the company’s records and related documents:

  1. On July 1, 2018, a two-year insurance premium on equipment in the amount of $480 was paid and debited in full to Prepaid Insurance on that date. Coverage began on July 1.
  2. At the end of 2018, the unadjusted balance in the Supplies account was $1,000. A physical count of supplies on December 31, 2018, indicated supplies costing $250 were still on hand.
  3. On December 31, 2018, YY’s Garage completed repairs on one of Brokeback’s trucks at a cost of $750. The amount is not yet recorded. It will be paid during January 2019.
  4. On December 31, 2018, the company completed a contract for an out-of-state company for $7,700 payable by the customer within 30 days. No cash has been collected and no journal entry has been made for this transaction.
  5. On July 1, 2018, the company purchased a new hauling van. Depreciation for July–December 2018, estimated to total $2,500, has not been recorded.
  6. As of December 31, the company owes interest of $450 on a bank loan taken out on October 1, 2018. The interest will be paid when the loan is repaid on September 30, 2019. No interest has been recorded yet.
  7. Assume the income after the preceding adjustments but before income taxes was $25,000. The company’s federal income tax rate is 30%. Compute and record income tax expense.



Required:

  1. Give the adjusting journal entry required for each item at December 31, 2018.
  2. If adjustments were not made each period, the financial results could be materially misstated. Determine the amount by which Brokeback's net income would have been understated, or overstated, had the adjustments in requirement 1 not been made.
Answer #1
1
Debit Credit
a Insurance expense 120 =480*6/24
      Prepaid Insurance 120
b Supplies expense 750 =1000-250
     Supplies 750
c Repairs and Maintenance expense 750
      Accrued Liabilities/Accounts Payable 750
d Accounts Receivable 7700
      Service Revenue 7700
e Depreciation expense 2500
      Accumulated depreciation 2500
f Interest expense 450
      Interest payable 450
g Income tax expense 7500 =25000*30%
    Income tax payable 7500
2
Brokeback's net income would be overstated by 4370 =(120+750+750+2500+450+7500)-7700